OpenStudy (anonymous):

Each year terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid $130. This year, he paid $169. What was the percent increase in the cost of his plane ticket.? show your work. guys plzz. Help me I give medals

OpenStudy (help!!!!):

\[\frac{ change }{ originial}=\frac{ x }{ 100 }\]

OpenStudy (yanasidlinskiy):

169 - 130 = $39 (39/130) * 100 = 30% increase It's like a proportion like ^

OpenStudy (anonymous):

Thanks you guys

OpenStudy (yanasidlinskiy):

Welcome:)!!!!!

OpenStudy (anonymous):

Hey yana